# Kartik Narayan, Vibashan VS, Vishal M. Patel

### v2

@@ Abstract
1 Introduction
Recent progress in Large Language Models (LLMs) [2, 7, 102] have showcased their impressive ability to understand, reason, and generate text across diverse open-ended tasks. Building on these advancements, multimodal large language models (MLLMs) [51, 138, 59, 118, 52, 66] have rapidly emerged, using LLMs as a brain to process multimodal inputs including images, videos, and audio. These models [49, 12, 1, 56, 27] demonstrate remarkable performance and excel in challenging visual-question-answering (VQA) benchmarks such as MMMU [122], MME [120], MM-Bench [63], SEED-Bench [48], MMAU [119], MMIE [110], and MathVista [65]. These benchmarks evaluate MLLMs in various domains, including college-level subject knowledge, mathematical reasoning, commonsense reasoning, planning, coding, diagram and spatio-temporal understanding. Despite their progress, prominent MLLMs such as GPT-4o [31] and GeminiPro 1.5 [99] struggle with basic face understanding tasks, as illustrated in Figure. 1.
Before exploring MLLMs for face understanding capabilities, it’s essential to address the fundamental question: “Why should MLLMs be proficient in face understanding?” MLLMs are increasingly deployed as central processors in various advanced applications, including virtual-reality headsets [39], embodied AI [30, 21], driving safety [32, 92], authentication [19], human-computer interaction [103], and sports analysis [109]. In these applications, accurate face understanding is crucial, as face images appear frequently and require accurate face understanding for appropriate responses. However, the face understanding capabilities of existing MLLMs are limited; they often fail to answer basic questions such as “What is the expression of the person in this image?” or “What is the age of the person in the face image?” These shortcomings indicate a significant scope for improvement. Recent works, such as EMO-LLaMA [112], develop an instruction-tuning set for supervised fine-tuning to enhance MLLMs’ capabilities in understanding facial expressions. Face-MLLM [94] proposed a three-stage training pipeline to equip MLLMs’ with face perception capabilities. Nonetheless, the research community currently lacks a standardized benchmark to quantify and compare MLLMs’ performance in face understanding. We believe that a comprehensive benchmark incorporating various aspects of face understanding is a crucial and essential foundational step for monitoring progress and advancing MLLMs’ performance in this domain. A more detailed discussion, addressing additional questions that frame our motivation, is provided in Appendix A.
@@ 1 Introduction
  • Extensive Evaluation: We evaluate $26$ open-source MLLMs and two proprietary models, GPT-4o and GeminiPro1.5, which achieve accuracies of $50.24$% and $54.40$%, respectively, highlighting the significant challenge posed by FaceXBench and the substantial room for improvement.
@@
2 Related Work
Face Understanding: Face understanding concerns the analysis and interpretation of human facial features, encompassing tasks such as age estimation [93, 38, 46, 10], gender prediction [40, 5, 57], race estimation [3, 4], face recognition [74, 104, 20], face anti-spoofing [116, 61, 72], deepfake detection [131, 100, 73], facial attributes prediction [69, 139, 126], facial expression recognition [105, 44, 123], headpose estimation [125, 26, 83], crowd counting [124, 80, 79] and face parsing [133, 86, 76, 97]. Early methods focused on developing task-specific models for each facial understanding task, achieving promising results. With the advent of transformers, works such as FaceXFormer [75], Q-Face [95] and Faceptor [77] aim to unify multiple tasks within a single architecture to improve generalization performance. The recent rise in MLLMs’ enabled new avenue of research in face understanding by integrating information across text, visual and other modalities. Recent works [15, 115, 132, 89, 112, 94], leverage the reasoning and zero-shot capabilities of MLLMs’ to approach traditional face-related tasks. However, the field still lacks a standardized benchmark to monitor and regulate the development of these models. Our proposed work addresses this gap by introducing a comprehensive benchmark, FaceXBench, which covers a diverse range of face understanding tasks.
@@ 2 Related Work
3 The FaceXBench Benchmark
3.1 Overview of FaceXBench
@@
Benchmark Statistics. FaceXBench consists of $5,000$ questions, divided into $6$ categories covering $14$ tasks, and is derived from $25$ public datasets along with one proposed dataset (i.e. FaceXAPI). There are $2,750$ questions with multiple images, $2,150$ single-image questions, and $100$ text-only questions. The questions are generated using $757$ unique question templates and contain $10,441$ unique images. The correct answer options are approximately equally distributed between A, B, C and D to avoid bias. They key statistics of the benchmark are summarized in Table 1.
Table 1: Key statistics of questions in FaceXBench.
@@
Questions with multiple images 2750 (55%)
Questions with single image 2150 (43%)
@@
Total images in all questions 11266
Unique number of images 10441
Unique question templates 757
@@
Average question length 64.34
Average option length 11.04
@@
Frequency of A as correct option 1278 (25.56%)
Frequency of B as correct option 1332 (26.64%)
Frequency of C as correct option 1189 (23.78%)
Frequency of D as correct option 1201 (24.02%)
3.2 Data Collection
@@
Our data collection pipeline consists of three steps. Step 1. In the first step, we iterate through the identified tasks in each category and select the datasets corresponding to each task. We collect the test sets of existing standard datasets for each task to avoid data leakage. We also created a new dataset, FaceXAPI, specifically for the tool-retrieval task. In total, our collection includes $25$ public datasets along with this newly proposed dataset. Step 2. In this step, we manually created question templates for each task involving single and multiple images. While creating these templates, we carefully framed questions to encourage the model to reason, compare, and think critically to find the correct answer. We included a mix of easy and hard questions to maintain a balanced distribution. For example, a hard question is, “<image1><image2><image3> How many images show a person in the age range of $30$ to $39$?,” while an easy question is, “<image1> What is the age range of the person shown in the image?”. We prompt GPT-4o with the manually curated question templates as in-context examples and generate additional templates that address similar questions but with varied phrasing, perspective, and/or complexity. For example, “<image1><image2> Which among these two appears to be older?”. We manually filter the question templates for each task to ensure diversity and task relevance. Additional details about the question templates and their generation can be found in Appendix D.1. Step 3. The final step involves generating answer options. Each question includes four options, with one correct answer. We ensure that distractor options are close to the correct answer but distinct enough to require critical thinking and reasoning when selecting the correct response. For example, for the question “<image1> What is the age range of the person shown in the image?”, the options designed are (A) $20$ to $29$, (B) $10$ to $19$ (C) $30$ to $39$ and (D) $40$ to $49$, with the correct answer being (C) $20$ to $29$. The thoughtfully chosen distractor options close to the actual answer, encourages the model to reason, compare and carefully select the correct choice. This approach raises the difficulty for the model, making the questions more challenging. To validate this, we conduct a mini experiment using LLaVA-OV for age estimation on FairFace dataset questions. The model achieved $88$% accuracy with randomly chosen options, compared to $53$% when strategically designed options were used. This result highlights the importance of identifying effective distractor options. Further details on generating distractor options for each task are provided in Appendix D.2. Overall, we observed that the model’s performance significantly varies depending on the complexity of the questions and distractor options.
@@
3.3 FaceXAPI
We believe that face understanding is an application domain that can be better addressed by equipping MLLMs with tools rather than relying solely on supervised fine-tuning. We provide a detailed discussion, supported by experimental validation, in the Discussion and Future Directions section (Section 6). To test MLLMs’ ability to select the correct sequence of APIs and function calls for successful task completion, we created FaceXAPI, a dataset of $100$ text-only questions referencing 13 APIs and 32 function calls. The questions were designed to ensure diversity in scenarios, reflecting a broad range of real-world applications and requiring a sequence of $3$ to $5$ API calls to solve. The dataset includes a total of 88 unique function call sequences. The questions were generated using GPT-4o, followed by manual filtering. A sample from the dataset is presented in Figure 3, with the complete prompt provided in Appendix D.3.
3.4 Quality Control
@@ 4 Experiments
In this section, we detail the experiments conducted to benchmark and analyze various MLLMs on face understanding. We evaluate $2$ proprietary and $26$ open-source models, listed in Section 4.1. For a fair comparison, all models are evaluated in a zero-shot setting using the same base prompt. We analyze selected MLLMs under in-context and chain-of-thought settings. The evaluation settings and results are described in Sections 4.2 and 5, respectively.
4.1 Models
@@
The $2$ proprietary models used are GPT-4o [31] and GeminiPro 1.5 [99]. We divide the $26$ open-source models into three major categories based on parameter size: (a) Open Source MLLMs (<4B parameters): PaliGemma [9], LLaVA-OneVision-0.5b-OV [49], and VILA 1.5-3b [56]; (b) Open Source MLLMs (4B-13B parameters): Chameleon-7b [98], Eagle-X4-8B-Plus [88], Idefics2-8b [42], Idefics-9b-Instruct [41], LLaVA-v1.5-7b [59], Monkey-Chat [55], MiniCPM-Llama3-v2.5 [117], LLaVA-OneVision-7b-SI [49], LLaVA-NeXT-Interleave-7b [50], Mantis-SIGLIP-8b [33], Phi-3.5-Vision [1], LLaVA-OneVision-7b-OV, Qwen2-VL-7b-Instruct [106], and InternVL2-8b [12]; (c) Open Source MLLMs (>13B parameters): CogVLM2-19b [27], Idefics-80b-Instruct [41], LLaVA-v1.5-13b [59], VILA 1.5-13b [56], InternVL-Chat-v1.5 [12], VILA 1.5-40b [56], LLaVA-OneVision-72b-OV [49], Qwen2-VL-72b-Instruct [106], and InternVL2-76b [12]. In the Appendix E.1, we provide detailed information regarding the architecture and the parameter size for all open-source MLLMs’ evaluated, along with additional results under different settings.
4.2 Evaluation Settings
@@
We evaluate the models under three settings: (a) zero-shot, (b) in-context task description, and (c) chain-of-thought prompting. In the zero-shot setting, we pass the input with only the base prompt. In the in-context task description setting, we prepend the prompt with a brief description of the specific task required by the question. In the chain-of-thought setting, we prompt the model to reason step-by-step before selecting the correct option. The task-specific prepended text and additional details of the evaluation settings are provided in the Appendix D.6.
@@
The proposed benchmark consists of multiple-choice questions (MCQs), which standardize the evaluation across different models. We empirically construct a diverse set of regular expressions and employ a three-step evaluation strategy to extract the correct option in cases where intermediate reasoning or calculations are present in the model’s output. Our evaluation strategy is as follows: First, we use a regular expression to match the correct option (A, B, C, or D) at the beginning of the model’s prediction. If this attempt fails, we then search for the correct option within the entire prediction text. As a final fallback, we compare parts of the predicted output with the option values to find a match. If none of these steps succeed, we categorize the prediction as incorrect. The complete function is provided in Appendix D.4. Alternatively, we could implement a random-choice or frequent-choice strategy instead of labeling responses as incorrect. In the random-choice approach, a random option is selected as the prediction, while in the frequent-choice approach, we select the most common correct option in the dataset. However, these strategies can lead to misleading results, especially when the model fails to respond due to content moderation measures or poor instruction-following.
5 Results
@@
Open source MLLMs ($\mathbf{4}\text{B}$ $\mathbf{13}\text{B}$ parameters)
@@
Table 2: Results of different models on the FaceXBench. We categorize the open-source models in three categories based on parameter size: (a) Open source MLLMs (<4B parameters), (b) Open source MLLMs (4B-13B parameters), (c) Open source MLLMs (>13B parameters). Additionally, we evaluate (d) proprietary models. The best-performing model in each category is highlighted in bold.
@@
The performance of various models on FaceXBench is shown in Table 2. The results emphasize the challenging nature of the benchmark, with no model achieving more than $60$% accuracy. We observe that Qwen2-VL-72b-Instruct [106] achieves the best overall performance of 57.86%. InternVL2-76b [12], GeminiPro 1.5 [99], Qwen2-VL-72b-Instruct [106], LLaVA-OneVision-72b-OV [49], Qwen2-VL-72b-Instruct [106] and GeminiPro 1.5 [99] achieve the highest performances in the categories of Bias & Fairness, Face Recognition, Face Authentication, Face Analysis, Face Localization and Face Tools Use with accuracies of $69.53$%, $70.00$%, $41.14$%, $63.25$%, $55.45$% and $57.00$%, respectively. Human performance at $70.28$% and SOTA vision models at $84.50$%, serve as the upper bounds for the benchmark. The significant gap between the current state-of-the-art MLLM and this upper bound highlights the substantial room for improvement.
Results across different evaluation categories. The performance of various models on FaceXBench, as shown in Table 2 highlights the limited face understanding capabilities of current models. Specifically, they perform poorly on tasks such as face authentication and face localization, which require fine-grained facial feature extraction and spatial understanding of facial structure. The low performance in the “Bias & Fairness” category suggests that existing MLLMs exhibit biases toward certain age, gender, and racial groups, which need to be mitigated prior to deployment. Surprisingly, GPT-4o performs poorly in the “Bias & Fairness category”, but on examination, it was found that the model often chose not to answer due to its safety alignment. The top-performing models achieve around $70$% performance in face recognition tasks; however, their performance drops significantly on low-resolution face recognition. This decline can be attributed to training data predominantly comprising of high-resolution images, which limits models’ effectiveness on low-resolution inputs. Some models may have been trained on image-text datasets with face images, such as FFHQ-Text [137], CelebA-Dialog [34], LAION-Face [136], FaceCaption-15M [18], as part of large-scale pretraining. Although these datasets often contain attribute and expression information in text, the models still perform poorly in the “Face Analysis” section, indicating substantial room for improvement.
Open Source vs Closed Source Models. We observe that open-source models, such as InternVL2 and Qwen2-VL, outperform proprietary models like GPT-4o and GeminiPro 1.5, achieving accuracies of $57.80$% and $57.86$%, compared to $50.50$% and $56.96$%, respectively. Generally, MLLMs have shown a trend where closed-source models outperform open-source ones. However, in the sensitive domain of face analysis, proprietary models are safety-aligned before deployment. The relatively poor performance of proprietary models in “Bias & Fairness” and “Face Analysis” is primarily due to content moderation constraints. Notably, GeminiPro 1.5 demonstrates comparatively better performance than other models on “Face Tools Use”, showcasing its ability to leverage specialized tools for complex scenarios involving multiple face-related tasks.
Performance across various tasks. To gauge the difficulty of various tasks, we plot the performance of the top-5 models ($4$B-$13$B parameters) across each task, as shown in Figure 4(a). A detailed table showing the performance of all models across all tasks is provided in Appendix E.2. We observe that, on average, models struggle with tasks such as crowd counting, deepfake detection, head pose estimation, and low-resolution face recognition, highlighting areas where MLLMs’ need improvement. In contrast, gender prediction is one of the easier tasks, with an average performance of approximately $80$%. Additionally, we plot the average performance of these top-5 models on multiple-image and single-image question. Figure 4(c) shows that models generally perform poorly on questions with multiple images as inputs compared to single-image questions. This is primarily because the models are required to process more visual information and compare and contrast facial features across multiple images, making the questions more challenging. Furthermore, models such as Qwen2-VL [106] and LLaVA-OneVision [49], which use dynamic resolution to handle arbitrary image resolutions, perform better than models on segmentation task. This mapping of images to a dynamic number of visual tokens more closely resembles human-like visual processing and leads to improved performance in face-understanding tasks.
Effect of LLM and it’s size on performance. To analyze the impact of LLM on model performance, we plot a performance curve for different models that share the same SigLIP SO400M/14@384 vision encoder, but differ in their LLM backbone. From Figure 4(b), we observe that the performance improves with an increase in LLM size. Furthermore, when comparing LLMs of different sizes within the same family, such as Qwen2, we see that the shape of the curve remains consistent, with performance values translating upward across all dimensions. This indicates that as the LLM size increases, the model’s capabilities improve proportionately across various dimensions, maintaining a similar pattern of performance enhancement.
Performance under different evaluation settings. We evaluate three selected models under the in-context task description and chain-of-thought settings, with results summarized in Table. 3. In the in-context setting, we observe some performance improvements in face authentication and face tools use. However, overall performance drops, indicating that the models struggle to utilize in-context information effectively. In the chain-of-thought setting, we observe a substantial decline in performance, suggesting that, although these models exhibit reasoning capabilities, they do not transfer effectively to face understanding.
@@
Figure 4: (a) Performance of top-5 models ($4$B-$13$B parameters) across various tasks. (b) Effect of LLM and it’s size on model performance. (c) Average performance of the top-5 models ($4$B-$13$B parameters) on multiple-image and single-image questions.
6 Discussion and Future Directions
In this section, we discuss possible future directions to improve the face understanding capabilities of MLLMs. The MLLM community has developed numerous supervised fine-tuning (SFT) datasets, such as COCO Caption [11], ScienceQA [84], Vision-FLAN [113], ChartQA [67], FigureQA [36], Geometry3K [24], MAVIS MCollect [127], MATHQA [6], TextOCR [91], OCR-VQA [68], and MagpiePro [114]. These supervised fine-tuning sets target various skills, including document and chart understanding, mathematics, fine-grained perception, grounding, reasoning, and general OCR. Developing these general skills helps models perform better on existing tasks and enhances their reasoning and visual processing capabilities for improved zero-shot task performance. However, to answer the question, “Will supervised fine-tuning improve the performance of MLLMs on face understanding?” we conducted multiple experiments by finetuning LLaVA1.5 [58] on a random $70$k subset of the FaceCaption15M [18], which comprises image-text pairs containing attribute, age, and gender information. We fine-tuned the vision projector and the LLM backbone using LoRA [28] and summarized the results for various data compositions in Table 3(c). Our findings reveal that naive fine-tuning on MLLM with only Face SFT data results in poor performance (Table 3(c), row 1), as models tend to lose their general reasoning and perception capabilities. We incorporated the complete LLaVA SFT dataset ($665$k samples) alongside the Face SFT, and the model achieved a score of 35.18 (Table 3(c), row 2). We randomly sampled $200$k examples from the full $665$k LLaVA 1.5 dataset, combining them with the 70k Face SFT data for fine-tuning, resulting in an improved score of $35.24$ (Table 3(c), row 3). This experiment demonstrates the benefit of integrating FaceSFT data in the right proportion with existing reasoning and instruction-tuning datasets.
We explored an alternative research direction to improve MLLMs’ performance on face understanding through the use of specialized tools. To investigate the benefits of tool use, we selected specific datasets and converted predictions from state-of-the-art models into text, which we then provided as context to the MLLM. As shown in Table 3(b), this approach resulted in a significant performance boost. We believe that equipping MLLMs with agentic behavior through specialized tools is a promising way to enhance their face understanding capabilities.
For future work: we advocate the development of diverse supervised fine-tuning sets covering various aspects of face understanding and training MLLMs to leverage specialized tools. Following this direction, we believe that future MLLMs will be capable of advanced face understanding, with FaceXBench serving as a critical resource for monitoring progress by benchmarking models across multiple dimensions of face understanding.

### v3

# on Face Understanding

@@ Abstract
Index Terms:
Multimodal LLMs, Benchmark, Face Understanding
I Introduction
Recent advances in large language models (LLMs)[2, 7, 102] have demonstrated strong capabilities in understanding, reasoning, and generating text across diverse open-ended tasks. Building on this progress, multimodal LLMs (MLLMs)[50, 138, 58, 118, 51, 65] have emerged, leveraging LLMs to process multimodal inputs such as images, videos, and audio. These models [46, 12, 1, 55, 27] excel in visual-question-answering (VQA) benchmarks like MMMU [122], MME [120], MM-Bench [61], SEED-Bench [47], MMAU [119], MMIE [110], and MathVista [64], which test a range of skills including subject knowledge, mathematical and commonsense reasoning, planning, coding, and spatio-temporal understanding. Despite this progress, leading MLLMs such as GPT-4o [31] and GeminiPro 1.5 [99] still struggle with basic face understanding tasks (see Figure 1).
Before exploring MLLMs for face understanding capabilities, it’s essential to address the fundamental question: “Why should MLLMs be proficient in face understanding?” MLLMs are increasingly deployed as central processors in various advanced applications, including virtual-reality headsets [39], embodied AI [30, 21], driving safety [32, 92], authentication [19], human-computer interaction [103], and sports analysis [109]. In these applications, accurate face understanding is crucial, as face images appear frequently and require accurate face understanding for appropriate responses. However, the face understanding capabilities of existing MLLMs are limited; they often fail to answer basic questions such as “What is the expression of the person in this image?” or “What is the age of the person in the face image?” These shortcomings indicate a significant scope for improvement. Recent works, such as EMO-LLaMA [112], develop an instruction-tuning set for supervised fine-tuning to enhance MLLMs’ capabilities in understanding facial expressions. Face-MLLM [93] proposed a three-stage training pipeline to equip MLLMs’ with face perception capabilities. Nonetheless, the research community currently lacks a standardized benchmark to quantify and compare MLLMs’ performance in face understanding. We believe that a comprehensive benchmark incorporating various aspects of face understanding is a crucial and essential foundational step for monitoring progress and advancing MLLMs’ performance in this domain. A more detailed discussion, addressing additional questions that frame our motivation, are provided in the Appendix.
@@ I Introduction
  • Extensive Evaluation: We evaluate $26$ open-source MLLMs and two proprietary models, GPT-4o and GeminiPro1.5, which achieve accuracies of 50.2450.24% and 54.4054.40%, respectively, highlighting the significant challenge posed by FaceXBench and the substantial room for improvement.
@@
II Related Work
Face Understanding: Face understanding concerns the analysis and interpretation of human facial features, encompassing tasks such as age estimation [95, 38, 45, 10], gender prediction [40, 5, 56], race estimation [3, 4], face recognition [73, 105, 20], face anti-spoofing [116, 62, 74], deepfake detection [131, 100, 72, 71], facial attributes prediction [68, 139, 126], facial expression recognition [104, 43, 123], headpose estimation [124, 26, 83], crowd counting [125, 80, 79] and face parsing [133, 86, 76, 97]. Early methods focused on developing task-specific models for each facial understanding task, achieving promising results. With the advent of transformers, works such as FaceXFormer [75], Q-Face [94] and Faceptor [77] aim to unify multiple tasks within a single architecture to improve generalization performance. The recent rise in MLLMs’ enabled new avenue of research in face understanding by integrating information across text, visual and other modalities. Recent works [15, 114, 132, 89, 112, 93], leverage the reasoning and zero-shot capabilities of MLLMs’ to approach traditional face-related tasks. However, the field still lacks a standardized benchmark to monitor and regulate the development of these models. Our proposed work addresses this gap by introducing a comprehensive benchmark, FaceXBench, which covers a diverse range of face understanding tasks.
@@ II Related Work
III The FaceXBench Benchmark
III-A Overview of FaceXBench
@@
TABLE I: Key statistics of questions in FaceXBench.
@@
Questions with multiple images 2900 (58%)
Questions with single image 2000 (40%)
@@
Total images in all questions 11694
Unique number of images 10735
Unique question templates 573
@@
Average question length 67.13
Average option length 11.55
@@
Frequency of A as correct option 1256 (25.12%)
Frequency of B as correct option 1266 (25.32%)
Frequency of C as correct option 1246 (24.92%)
Frequency of D as correct option 1232 (24.64%)
Benchmark Statistics. FaceXBench consists of $5,000$ questions, divided into $6$ categories covering $14$ tasks, and is derived from $25$ public datasets along with one proposed dataset (i.e. FaceXAPI). There are $2,900$ questions with multiple images, $2,000$ single-image questions, and $100$ text-only questions. The questions are generated using $573$ unique question templates and contain $10,735$ unique images. The correct answer options are approximately equally distributed between A, B, C and D to avoid bias. The key statistics of the benchmark are summarized in Table I.
III-B Data Collection
@@
Our data collection pipeline consists of three steps. Step 1. In the first step, we iterate through the identified tasks in each category and select the datasets corresponding to each task. We collect the test sets of existing standard datasets for each task to avoid data leakage. We also created a new dataset, FaceXAPI, specifically for the tool-retrieval task. In total, our collection includes $25$ public datasets along with this newly proposed dataset. Step 2. In this step, we manually created question templates for each task involving single and multiple images. While creating these templates, we carefully framed questions to encourage the model to reason, compare, and think critically to find the correct answer. We included a mix of easy and hard questions to maintain a balanced distribution. For example, a hard question is, “<image1><image2><image3> How many images show a person in the age range of $30$ to 3939?,” while an easy question is, “<image1> What is the age range of the person shown in the image?”. We prompt GPT-4o with the manually curated question templates as in-context examples and generate additional templates that address similar questions but with varied phrasing, perspective, and/or complexity. For example, “<image1><image2> Which among these two appears to be older?”. We manually filter the question templates for each task to ensure diversity and task relevance. Additional details about the question templates and their generation can be found in the supplementary material. Step 3. The final step involves generating answer options. Each question includes four options, with one correct answer. We ensure that distractor options are close to the correct answer but distinct enough to require critical thinking and reasoning when selecting the correct response. For example, for the question “<image1> What is the age range of the person shown in the image?”, the options designed are (A) $20$ to $29$, (B) $10$ to $19$ (C) $30$ to $39$ and (D) $40$ to $49$, with the correct answer being (A) $20$ to $29$. The thoughtfully chosen distractor options close to the actual answer, encourages the model to reason, compare and carefully select the correct choice. This approach raises the difficulty for the model, making the questions more challenging. To validate this, we conduct a mini experiment using LLaVA-OV for age estimation on FairFace dataset questions. The model achieved 8888% accuracy with randomly chosen options, compared to 5353% with strategically designed options. This result highlights the importance of identifying effective distractor options. Further details on generating distractor options for each task are provided in the supplementary material. Overall, we observed that the model’s performance significantly varies with the complexity of the questions and distractor options.
@@
III-C FaceXAPI
We believe that face understanding is an application domain that can be better addressed by equipping MLLMs with tools rather than relying solely on supervised fine-tuning. We provide a detailed discussion, supported by experimental validation, in the Discussion and Future Directions section (Section VI). To test MLLMs’ ability to select the correct sequence of APIs and function calls for successful task completion, we created FaceXAPI, a dataset of $100$ text-only questions referencing 13 APIs and 32 function calls. The questions were designed to ensure diversity in scenarios, reflecting a broad range of real-world applications and requiring a sequence of $3$ to $5$ API calls to solve. The dataset includes a total of 88 unique function call sequences. The questions were generated using GPT-4o, followed by manual filtering. A sample from the dataset is presented in Figure 2, with the complete prompt provided in the supplementary material.
III-D Quality Control
@@ IV Experiments
In this section, we detail the experiments conducted to benchmark and analyze various MLLMs on face understanding. We evaluate $2$ proprietary and $26$ open-source models, listed in Section IV-A. For a fair comparison, all models are evaluated in a zero-shot setting using the same base prompt. We analyze selected MLLMs under in-context and chain-of-thought settings. The evaluation settings and results are described in Sections IV-B and V, respectively.
IV-A Models
@@
The $2$ proprietary models used are GPT-4o [31] and GeminiPro 1.5 [99]. We divide the $26$ open-source models into three major categories based on parameter size: (a) Open Source MLLMs (<4B parameters): PaliGemma [9], LLaVA-OneVision-0.5b-OV [46], and VILA 1.5-3b [55]; (b) Open Source MLLMs (4B-13B parameters): Chameleon-7b [98], Eagle-X4-8B-Plus [88], Idefics2-8b [laurençon2024matters], Idefics-9b-Instruct [41], LLaVA-v1.5-7b [58], Monkey-Chat [54], MiniCPM-Llama3-v2.5 [117], LLaVA-OneVision-7b-SI [46], LLaVA-NeXT-Interleave-7b [49], Mantis-SIGLIP-8b [33], Phi-3.5-Vision [1], LLaVA-OneVision-7b-OV, Qwen2-VL-7b-Instruct [106], and InternVL2-8b [12]; (c) Open Source MLLMs (>13B parameters): CogVLM2-19b [27], Idefics-80b-Instruct [41], LLaVA-v1.5-13b [58], VILA 1.5-13b [55], InternVL-Chat-v1.5 [12], VILA 1.5-40b [55], LLaVA-OneVision-72b-OV [46], Qwen2-VL-72b-Instruct [106], and InternVL2-76b [12]. In the supplementary material, we provide detailed information regarding the architecture and the parameter size for all open-source MLLMs’ evaluated, along with additional results under different settings.
IV-B Evaluation Settings
@@
We evaluate the models under three settings: (a) zero-shot, (b) in-context task description, and (c) chain-of-thought prompting. In the zero-shot setting, we pass the input with only the base prompt. In the in-context task description setting, we prepend the prompt with a brief description of the specific task required by the question. In the chain-of-thought setting, we prompt the model to reason step-by-step before selecting the correct option. The task-specific prepended text and additional details of the evaluation settings are provided in the supplementary material.
@@
The proposed benchmark consists of multiple-choice questions (MCQs), which standardize the evaluation across different models. We empirically construct a diverse set of regular expressions and employ a three-step evaluation strategy to extract the correct option in cases where intermediate reasoning or calculations are present in the model’s output. Our evaluation strategy is as follows: First, we use a regular expression to match the correct option (A, B, C, or D) at the beginning of the model’s prediction. If this attempt fails, we then search for the correct option within the entire prediction text. As a final fallback, we compare parts of the predicted output with the option values to find a match. If none of these steps succeed, we categorize the prediction as incorrect. The complete function is provided in the supplementary material. Alternatively, we could implement a random-choice or frequent-choice strategy instead of labeling responses as incorrect. In the random-choice approach, a random option is selected as the prediction, while in the frequent-choice approach, we select the most common correct option in the dataset. However, these strategies can lead to misleading results, especially when the model fails to respond due to content moderation measures or poor instruction-following.
IV-C Implementation Details
  • VLMEvalKit: We utilized the official open-source implementation shared and maintained on GitHub, available at https://github.com/open-compass/VLMEvalKit.
V Results
@@
Open source MLLMs ($\mathbf{4}\text{B}$ - $\mathbf{13}\text{B}$ parameters)
@@
TABLE II: Results of different models on the FaceXBench. We categorize the open-source models in three categories based on parameter size: (a) Open source MLLMs (<4B parameters), (b) Open source MLLMs (4B-13B parameters), (c) Open source MLLMs (>13B parameters). Additionally, we evaluate (d) proprietary models. The best-performing model in each category is highlighted in bold.
@@
The performance of various models on FaceXBench is shown in Table II. The results emphasize the challenging nature of the benchmark, with no model achieving more than 6060% accuracy. We observe that Qwen2-VL-72b-Instruct [106] achieves the best overall performance of 57.86%. InternVL2-76b [12], GeminiPro 1.5 [99], Qwen2-VL-72b-Instruct [106], LLaVA-OneVision-72b-OV [46], Qwen2-VL-72b-Instruct [106] and GeminiPro 1.5 [99] achieve the highest performances in the categories of Bias & Fairness, Face Recognition, Face Authentication, Face Analysis, Face Localization and Face Tools Use with accuracies of 69.5369.53%, 70.0070.00%, 41.1441.14%, 63.2563.25%, 55.4555.45% and 57.0057.00%, respectively. Human performance at 70.2870.28% and SOTA vision models at 84.5084.50%, serve as the upper bounds for the benchmark. The significant gap between the current state-of-the-art MLLM and this upper bound highlights the substantial room for improvement.
Models Overall
Expression
Recognition
Age
Estimation
Race
Estimation
Crowd
Counting
Celebrity
Identification
HR FR LR FR
Face
Anti-spoofing
Tools
Retrieval
Attributes
Prediction
Gender
Prediction
Headpose
Estimation
Face
Parsing
Deepfake
Detection
Open source MLLMs ($\mathbf{<4}\text{B}$ parameters)
Open source MLLMs ($\mathbf{4}\text{B}$ - $\mathbf{13}\text{B}$ parameters)
Open source MLLMs ($\mathbf{>13}\text{B}$ parameters)
Proprietary MLLMs
TABLE III: Performance of various models across all evaluation tasks of FaceXBench.
@@
Figure 4: (a) Performance of top-5 models (44B-1313B parameters) across various tasks. (b) Effect of LLM and it’s size on model performance. (c) Average performance of the top-5 models (44B-1313B parameters) on multiple-image and single-image questions.